Project Information for Quiz
You have accepted a project that asks that you build 20 birdhouses at $60 apiece in 4 weeks.
It is now the first day of week 2 and you have expended (spent) $200 of your budget, completed 4 birdhouses, and want to check your earned value as part of cost management.

Question 1

Not yet answered

Points out of 2.0

Flag question

Question text

What is the planned value (PV) at this point, in dollars?

Select one:

a. $60

b. $240

c. $300

d. $360

Clear my choice

Question 2

Not yet answered

Points out of 2.0

Flag question

Question text

What is the actual cost (AC)?

Select one:

a. $200

b. $240

c. $300

d. $1600

Clear my choice

Question 3

Not yet answered

Points out of 2.0

Flag question

Question text

What is the earned value (EV)?

Select one:

a. $200

b. $240

c. $300

d. $1600

Clear my choice

Question 4

Not yet answered

Points out of 2.0

Flag question

Question text

What is the cost variance (CV)?

Select one:

a. $40

b. $60

c. $120

d. $240

Clear my choice

Question 5

Not yet answered

Points out of 2.0

Flag question

Question text

Is your project on budget, over budget, or under budget?

Select one:

a. The project is right on budget at this point.

b. The project is over budget at this point.

c. The project under budget at this point.

Clear my choice
Answer #1

What is the planned value (PV) at this point, in dollars?

20 birdhouses in 4 weeks. Assuming linear implementation, the plan is to build 5 birdhouses each week. At 60 a piece this would come to Option C: $ 300

What is the actual cost (AC)?

Actual cost is =Actually spent budget = Option A: $ 200

What is the earned value (EV)?​​​​​​​

Earned value is % of work completed * Total Project value = (4/20)*(20*60)= 0.2*1200= Option B: $ 240

What is the cost variance (CV)?

Cost Variance = Earned Value - Actual Cost = 240-200 = Option A: $ 40

Is your project on budget, over budget, or under budget?

Option C: Project is Under budget